# MIT-CTP/5802

### v1

@@ 1 Introduction
The first modern ML applications in LHC physics were jet taggers [2, 3], which aimed to optimally analyze the substructure of jets to identify their partonic nature [4, 5] based on all available low-level information [4, 5, 6, 7, 8]. These jet taggers also triggered the question of how and whether to include Lorentz symmetry [9], leading to a set of Lorentz-equivariant jet taggers, currently under experimental study [10, 11, 12].***In physics, we typically use the term “covariant” rather than “equivariant”, but equivariant networks are an established subfield of ML, so we stick to this term. Targeting the ultimate tagging precision with limited training data, these equivariant taggers paved the way for a new phase of precision-machine learning for the LHC, where established ML methods have to be systematically enhanced for and by particle physics.
@@
Our goal is to provide a way to systematically exploit Lorentz equivariance in for different networks tasks defined on relativistic phase space. A Lorentz-equivariant transformer can provide the appropriate internal or latent representation of phase space data to regression, classification, and generative networks. L-GATr is based on the Geometric Algebra Transformer (GATr) [31, 32], designed for Euclidean translations, rotations, and reflections. We generalize it to L-GATr encoding Lorentz-equivariance into new network layers, including a maximally expressive linear map, attention, and layer normalization. This architecture was originally developed for an ML audience and applied to amplitude regression, top tagging, and event generation in Ref. [33]. In this physics-targeted study, we extend the amplitude regression part, improve the classification through pre-training and multi-class tagging, and deliver a competitive generative network. We also provide a comprehensive benchmarking with the state of the art for these three distinct LHC applications.
@@ 2.2 Constructing a Lorentz-Equivariant Architecture
$\text{Linear}(x)$ $vx+w$ $\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}\;\left(+\sum_{k=0}^{4}w_{k}\gamma^{5}\langle x\rangle_{k}\right)$
@@ 4 L-GATr for Jet Tagging
We present the results in Tab. 4, including reference metrics from Refs. [57, 58]. The L-GATr tagger achieves a significant improvement over the previous state-of-the-art, ParT and MIParT, in essentially all signal types. In a separate study, we have checked that the quality of L-GATr predictions steadily increases as we add more features to the training data. We also train L-GATr with subsets of the dataset to test its data efficiency. As we can see in the left panel in Fig. 4, L-GATr achieves a performance similar to the non-equivariant ParT and MIParT taggers even if trained with only 10% of all available jets.
@@ 5 L-GATr for Event Generation
$x$ L-GATr rejection sampling -32.75 $\pm$ 0.05 0.530 $\pm$ 0.017
@@ Appendix A Network and Training Details
Architecture 128 channels 5 layers 128 channels 4 layers 128 channels 8 heads 8 blocks 96 scalar ch. 96 multivector ch. 8 blocks 72 scalar ch. 8 multivector ch. 4 blocks 32 scalar ch. 32 multivector ch. 8 heads 4 blocks

### v2

@@ 1 Introduction
The first modern ML applications in LHC physics were jet taggers [2, 3], which aimed to optimally analyze the substructure of jets to identify their partonic nature [4, 5] based on all available low-level information [4, 5, 6, 7, 8]. These jet taggers also triggered the question of how and whether to include Lorentz symmetry [9], leading to a set of Lorentz-equivariant jet taggers, currently under experimental study [10, 11, 12, 13, 14].***In physics, we typically use the term “covariant” rather than “equivariant”, but equivariant networks are an established subfield of ML, so we stick to this term. Targeting the ultimate tagging precision with limited training data, these equivariant taggers paved the way for a new phase of precision-machine learning for the LHC, where established ML methods have to be systematically enhanced for and by particle physics.
@@
Our goal is to provide a way to systematically exploit Lorentz equivariance in different networks tasks defined on relativistic phase space. A Lorentz-equivariant transformer can provide the appropriate internal or latent representation of phase space data to regression, classification, and generative networks. L-GATr is based on the Geometric Algebra Transformer (GATr) [33, 34], designed for Euclidean translations, rotations, and reflections. We generalize it to L-GATr encoding Lorentz-equivariance into new network layers, including a maximally expressive linear map, attention, and layer normalization. This architecture was originally developed for an ML audience and applied to amplitude regression, top tagging, and event generation in Ref. [35]. In this physics-targeted study, we extend the amplitude regression part, improve the classification through pre-training and multi-class tagging, and deliver a competitive generative network. We also provide a comprehensive benchmarking with the state of the art for these three distinct LHC applications.
@@ 2.2 Constructing a Lorentz-Equivariant Architecture
$\text{Linear}(x)$ $vx+w$ $\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}$
@@ 4 L-GATr for Jet Tagging
We present the results in Tab. 4, including reference metrics from Refs. [59, 60]. The L-GATr tagger achieves a significant improvement over the previous state-of-the-art, ParT and MIParT, in essentially all signal types. In a separate study, we have checked that the quality of L-GATr predictions steadily increases as we add more features to the training data. We also train L-GATr with subsets of the dataset to test its data efficiency. As we can see in the left panel in Fig. 4 and Table 5, L-GATr achieves a performance similar to the non-equivariant ParT and MIParT taggers even if trained with only 10% of all available jets.
@@
ParticleNet (2 M) 0.828 0.9820 5540 1681 90 662 1654 4049 4673 260 215
ParticleNet (10 M) 0.837 0.9837 5848 2070 96 770 2350 5495 6803 307 253
ParticleNet (100 M) 0.844 0.9849 7634 2475 104 954 3339 10526 11173 347 283
ParT (2 M) 0.836 0.9834 5587 1982 93 761 1609 6061 4474 307 236
ParT (10 M) 0.850 0.9860 8734 3040 110 1274 3257 12579 8969 431 324
ParT (100 M) 0.861 0.9877 10638 4149 123 1864 5479 32787 15873 543 402
MIParT (2 M) 0.837 0.9836 5495 1940 95 819 1778 6192 4515 311 242
MIParT (10 M) 0.850 0.9861 8000 3003 112 1281 3650 16529 9852 440 336
MIParT (100 M) 0.861 0.9878 10753 4202 123 1927 5450 31250 16807 542 402
L-GATr (2 M) 0.839 0.9842 6623 2294 99 981 1980 8097 4902 346 276
L-GATr (10 M) 0.859 0.9875 9804 3883 120 1791 4255 24691 13333 506 373
L-GATr (100 M) 0.866 0.9885 12987 4819 128 2311 6116 47619 20408 588 432
Table 5: Tagging accuracy, AUC, and background rejection $1/\epsilon_{B}$ on different sizes of the JetClass dataset [59]. Metrics from other models are taken from their published results [59, 60].
@@ 5 L-GATr for Event Generation
$x$ L-GATr rejection sampling -32.57 $\pm$ 0.05 0.530 $\pm$ 0.017
@@ Appendix A Network and Training Details
Architecture 128 channels 5 layers 128 channels 4 layers 128 channels 8 heads 8 blocks 96 scalar ch. 96 multivector ch. 8 blocks 72 scalar ch. 8 multivector ch. 4 blocks 32 scalar ch. 32 multivector ch. 8 heads 8 blocks